OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = x + 4 and passes through the point (−6, 3). y = 3/2x − 1 y = 3/2x + 3 y = −2/3x − 1 y = −2/3x + 3

OpenStudy (anonymous):

What would be the slope of a line perpendicular to the line \(y=x+4\)?

OpenStudy (anonymous):

You'll find all the information you need to answer this question here: http://www.purplemath.com/modules/slope2.htm

OpenStudy (anonymous):

i honestly don't know. i'm not understanding i was hoping someone could answer this step by step

OpenStudy (anonymous):

When you didn't respond, I thought you were away from your computer. A parallel line has the same slope. A perpendicular line has a slope that is the negative inverse.

OpenStudy (anonymous):

What is the slope of your original equation? \(y=x+4\)

OpenStudy (anonymous):

\(y=mx+b\) m = slope b = y-intercept